OpenStudy (anonymous):

Steve likes to entertain friends at parties with "wire tricks." Suppose he takes a piece of wire 68 inches long and cuts it into two pieces. Steve takes the first piece of wire and bends it into the shape of a perfect circle. He then proceeds to bend the second piece of wire into the shape of a perfect square. What should the lengths of the wires be so that the total area of the circle and square combined is as small as possible? (Round your answers to two decimal places.) I have attempted it several times and my answers add up to 68 but are wrong!

OpenStudy (ranga):

I am getting 29.91 inches for the circle wire and 38.09 inches for the square wire. If it is correct then I can type my steps here. If not I can just delete this answer.

OpenStudy (anonymous):

That is correct how did you find those numbers

OpenStudy (ranga):

The wire is 68 inches long. Assume C is the length of the wire cut for the circle. Then the length of the wire remaining for the square is (68 - C) C is the length of the circumference of a circle. So 2(pi)R = C, R = C/2pi Area of a circle = (pi)R^2 = (pi)C^2/(4pi^2) = C^2/(4pi) (cont'd)

OpenStudy (ranga):

The length of the wire for the square is (68 - C) The square has 4 equal sides adding up to (68 - C) So the side of the square must be (68-C)/4 The area of the square = (68-C)^2 / 16 Total area of circle + square = C^2/(4pi) + (68-C)^2 / 16

OpenStudy (anonymous):

then what do you do with the equations?

OpenStudy (ranga):

Let me put in values for pi and simplify the numbers Total Area = C^2/(4 x 3.1416) + (4624 + C^2 - 136C) / 16 = C^2 / 12.5664 + 289 + 0.0625C^2 - 8.5C = C^2 (1/12.5664 + 0.0625) - 8.5C + 289 = 0.1421C^2 - 8.5C + 289 = 0 To find the maximum/minimum, find the derivative with respect to C and set it to zero: dA/dC = 0.2842C - 8.5 = 0 C = 29.91 inches

OpenStudy (ranga):

So the wire length for the circle is 29.91 inches and the wire remaining for the square is (68 - 29.91) = 38.09 inches The second derivative is 0.2842 which is positive and therefore we have found a minimum Wire length for circle = 29.91 inches Wire length for square = 38.09 inches

OpenStudy (anonymous):

ok so then to find the total minimum area you plug the values back in

OpenStudy (ranga):

Correct. Total area A = 0.1421C^2 - 8.5C + 289 Put C = 29.91 and calculate A.

OpenStudy (ranga):

Total Area A = 161.89 square inches. If you want individual areas of the circle and the square, put the values in the individual equations for the area.

OpenStudy (anonymous):

I tried it and it did not provide the correct answer? Let me try the individual answers

OpenStudy (ranga):

The question you posted above only asks for the lengths.

OpenStudy (anonymous):

The second half is what is the combined minimal area.

OpenStudy (ranga):

It accepted the two lengths as correct but not 161.89 for the combined minimal area?

OpenStudy (anonymous):

Yes the lengths were right but 161.89 did not work

OpenStudy (ranga):

DO they tell you to what decimal place to round off the area?

OpenStudy (anonymous):

two decimals

OpenStudy (ranga):

Without any of the above formula you can recalculate the area of a circle formed by a wire 29.91 inches long and the area of a square formed by a wire of length 38.09 inches and you will get the area to be 161.89.

OpenStudy (anonymous):

Yea I have tried recalculating it and I get the same answer however it tells me its not correct. It is alright thank you for your help with finding the lengths

OpenStudy (ranga):

Try 161.87 It may be the second decimal that it is complaining about.

OpenStudy (anonymous):

It still didn't like it

OpenStudy (ranga):

Oh, well. But if you keep trying different second decimal it might eventually click. Doesn't make any sense. But anyway, you are welcome.